OpenStudy (anonymous):

solve the system of eq: x+ 2y = 6 -x/6 + y/3 =1

OpenStudy (anonymous):

x = 6-2y -(6-2y)/6 + y/3 = 1 -6+2y + 2y = 6 4y = 12 y=3 x=6-2*3 = 0